OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) find two positive real numbers whose product is 16 and whose difference is 1.

OpenStudy (anonymous):

xy = 16 x-y = 1 from second equation y = x-1 so x(x-1) = 16 x^2 - x - 16 = 0 x = [--1 +- sqrt((1 - 4*1*-6)] / 2 x = 4.53 so y = 3.53 the two numbers are 4.53 and 3.53

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

did you used quadratic formula?